# Robust Singing Voice Transcription Serves Synthesis

### v1

@@ Abstract
Note-level Automatic Singing Voice Transcription (AST) converts singing recordings into note sequences, facilitating the automatic annotation of singing datasets for Singing Voice Synthesis (SVS) applications. Current AST methods, however, struggle with accuracy and robustness when used for practical annotation. This paper presents ROSVOT, the first robust AST model that serves SVS, incorporating a multi-scale framework that effectively captures coarse-grained note information and ensures fine-grained frame-level segmentation, coupled with an attention-based pitch decoder for reliable pitch prediction. We also established a comprehensive annotation-and-training pipeline for SVS to test the model in real-world settings. Experimental findings reveal that ROSVOT achieves state-of-the-art transcription accuracy with either clean or noisy inputs. Moreover, when trained on enlarged, automatically annotated datasets, the SVS model outperforms its baseline, affirming the capability for practical application. Audio samples are available at https://rosvot.github.io.
@@ 3.4 Word Boundary Condition
Since in practice, the note annotation stage follows the phoneme annotation stage, word boundaries should already be obtained through forced alignment tools like MFA. We directly encode word boundaries as an additional condition to ensure word-note synchronization. However, to provide note-only support, we train an extra word boundary extractor $\text{E}_{\text{W}}$ to deal with scenarios like vocal tuning in music industries, where word alignment is unavailable. More details are listed in Appendix A.
@@ 4.1 Experimental Setup
Baselines  We compare ROSVOT, denoted as $\mathcal{M}$, with multiple baselines: 1) TONY Mauch et al. (2015), a automatic software with visualization; 2) VOCANO Hsu et al. (2021), retrained on the joint datasets; 3) MusicYOLO, retrained; 4) Yong et al. (2023), reproduced and retrained. We also compare the results of several variants of $\mathcal{M}$: 1) $\mathcal{M}$ (conformer), where the U-Net is dropped and the backbone is the Conformer alone; 2) $\mathcal{M}$ (conv), where the middle Conformer blocks are replaced by 8-layer convolution blocks; 3) $\mathcal{M}$ (w/o wbd), canceling word boundary condition; 4) $\mathcal{M}$ (w/o noise), which is identical to $\mathcal{M}$ but trained without noisy environment; 5) $\mathcal{M}$ (w/ $E_{\text{W}}$), meaning that the GT word boundaries are not available and need to be extracted from the extractor $\text{E}_{\text{W}}$.
@@ 4.4.2 SVS Results
We further investigate the performance as data size increases. While the AST model remains the same, we train $\mathcal{S}$ only using M4Singer as the baseline. Next, we gradually mix $\mathcal{D}_{1}$ and OpenSinger to expand the data size. To consume the largest datasets in the last row, we construct a large version of RMSSinger with 320-dimensional channels and a 6-layer decoder [^2] The dictionary of the text encoder is also merged with English phonemes for the following cross-lingual experiments, denoted as $\mathcal{S}$(large). The results are listed in Table 3, rows 8-11. A slight reduction in pitch accuracy can be observed when integrating diverse datasets, which may result from the inherent differences in dataset characteristics and annotation styles. However, the overall quality slightly improves, as the model has been exposed to a sufficient variety of pronunciation styles and singing patterns. This indicates that ROSVOT provides an opportunity for SVS models to scale up, which could be beneficial for large-scale singing voice pre-training or zero-shot SVS.
@@ 5 Conclusion
In this paper, we introduce ROSVOT, the first robust AST model that ultimately serves SVS. We leverage a multi-scale architecture to achieve a balance between coarse-grained note modeling and fine-grained segmentation. An attention-based decoder with dynamic weight is devised for pitch regression. Additionally, we establish a comprehensive pipeline for SVS training. Experimental results reveal that our model achieves the best performance under either clean or noisy environments. Annotating and incorporating larger datasets improves the SVS model’s performance, indicating the capability of practical annotation of ROSVOT.
6 Limitations and Potential Risks
@@ F.2 Additional Comparisons
  • Since each song in MIR-ST500 spans several minutes and needs to be source-separated first, it contains non-negligible un-voiced sections. We do not know whether Gu et al. (2023) segments the voiced parts according to some rules. Hence, we transcribe the whole song and compute the scores at once, resulting in a smaller proportion of positive samples of on/offsets due to longer silences, hereby influencing the performance. Another noteworthy point is that ROSVOT is able to load a 5-minute song into a 2080ti GPU and transcribe it at once, demonstrating considerable efficiency.

### v2

@@ Abstract
Note-level Automatic Singing Voice Transcription (AST) converts singing recordings into note sequences, facilitating the automatic annotation of singing datasets for Singing Voice Synthesis (SVS) applications. Current AST methods, however, struggle with accuracy and robustness when used for practical annotation. This paper presents ROSVOT, the first robust AST model that serves SVS, incorporating a multi-scale framework that effectively captures coarse-grained note information and ensures fine-grained frame-level segmentation, coupled with an attention-based pitch decoder for reliable pitch prediction. We also established a comprehensive annotation-and-training pipeline for SVS to test the model in real-world settings. Experimental findings reveal that ROSVOT achieves state-of-the-art transcription accuracy with either clean or noisy inputs. Moreover, when trained on enlarged, automatically annotated datasets, the SVS model outperforms its baseline, affirming the capability for practical application. Audio samples are available at https://rosvot.github.io. Codes can be found at https://github.com/RickyL-2000/ROSVOT.
@@ 3.4 Word Boundary Condition
Since in practice, the note annotation stage follows the phoneme annotation stage, word boundaries should already be obtained through forced alignment tools like MFA. We directly encode word boundaries as an additional condition to ensure word-note synchronization. Moreover, to provide note-only support, we train an extra word boundary extractor $\text{E}_{\text{W}}$ to deal with scenarios like vocal tuning in music industries, where word alignment is unavailable. More details are listed in Appendix A.
@@ 4.1 Experimental Setup
Baselines  We compare ROSVOT, denoted as $\mathcal{M}$, with multiple baselines: 1) TONY Mauch et al. (2015), an automatic software with visualization; 2) VOCANO Hsu et al. (2021), retrained on the joint datasets; 3) MusicYOLO, retrained; 4) Yong et al. (2023), reproduced and retrained. We also compare the results of several variants of $\mathcal{M}$: 1) $\mathcal{M}$ (conformer), where the U-Net is dropped and the backbone is the Conformer alone; 2) $\mathcal{M}$ (conv), where the middle Conformer blocks are replaced by 8-layer convolution blocks; 3) $\mathcal{M}$ (w/o wbd), canceling word boundary condition; 4) $\mathcal{M}$ (w/o noise), which is identical to $\mathcal{M}$ but trained without noisy environment; 5) $\mathcal{M}$ (w/ $E_{\text{W}}$), meaning that the GT word boundaries are not available and need to be extracted from the extractor $\text{E}_{\text{W}}$.
@@ 4.4.2 SVS Results
We further investigate the performance as data size increases. While the AST model remains the same, we train $\mathcal{S}$ only using M4Singer as the baseline. Next, we gradually mix $\mathcal{D}_{1}$ and OpenSinger to expand the data size. To consume the largest datasets in the last row, we construct a large version of RMSSinger with 320-dimensional channels and a 6-layer decoder [^2] The dictionary of the text encoder is also merged with English phonemes for the following cross-lingual experiments., denoted as $\mathcal{S}$(large). The results are listed in Table 3, rows 8-11. A slight reduction in pitch accuracy can be observed when integrating diverse datasets, which may result from the inherent differences in dataset characteristics and annotation styles. However, the overall quality slightly improves, as the model has been exposed to a sufficient variety of pronunciation styles and singing patterns. This indicates that ROSVOT provides an opportunity for SVS models to scale up, which could be beneficial for large-scale singing voice pre-training or zero-shot SVS.
@@ 5 Conclusion
In this paper, we introduced ROSVOT, the first robust AST model that ultimately serves SVS. We leveraged a multi-scale architecture to achieve a balance between coarse-grained note modeling and fine-grained segmentation. An attention-based decoder with dynamic weight was devised for pitch regression. Additionally, we established a comprehensive pipeline for SVS training. Experimental results revealed that our model achieved the best performance under either clean or noisy environments. Annotating and incorporating larger datasets improved the SVS model’s performance, indicating the capability of practical annotation of ROSVOT.
Limitations and Potential Risks
@@
Acknowledgement
This work was supported in part by the National Natural Science Foundation of China under Grant No.62222211.
@@ F.2 Additional Comparisons
  • Since each song in MIR-ST500 spans several minutes and needs to be source-separated first, it contains non-negligible un-voiced sections. We do not know whether Gu et al. (2023) segments the voiced parts according to some rules. Hence, we transcribe the whole song and compute the scores at once, resulting in a smaller proportion of positive samples of on/offsets due to longer silences, hereby influencing the performance. Another noteworthy point is that ROSVOT is able to load a 5-minute song into one single 2080ti GPU and transcribe it at once, demonstrating considerable efficiency.